What is the maximum flow rate of water in a smooth pipe of 8cm diameter if the flow is to be laminar


1
Expert's answer
2021-04-01T18:33:15-0400

Let's first find the average velocity of water in the pipe:


"R_e=\dfrac{\rho v_{avg}d}{\mu},"

here, "R_e=2000" is the Reynolds number, "\mu=10^{-3}" is the viscosity of water.

Then, we get:


"v_{avg}=\dfrac{R_e\mu}{\rho d}=\dfrac{2000\cdot10^{-3}\ Pa\cdot s}{1000\ \dfrac{kg}{m^3}\cdot0.08\ m}=0.025\ \dfrac{m}{s}."

The maximum velocity of water in the pipe will be twice the average velocity:


"v_{max}=2v_{avg}=2\cdot0.025\ \dfrac{m}{s}=0.05\ \dfrac{m}{s}=5.0\ \dfrac{cm}{s}."